please solve this as soon as possible

2. Let f:ZZ such that f(x) = 2x + 3. Prove that f one-to-one.
Answer #1

sor is de defined by f: 2-2 feu) = 2x +3. det f(x) = fy) for nyte. » 24+7=24+7 (using definition 24:27 opt). N- It ay Thus,